# Constraining pseudo-Diracness with astrophysical neutrino flavors

### v2

@@ V Conclusions
In this work, we have shown that using astrophysical neutrino flavor ratios, it is possible to distinguish between six possible cases of pseudo-Dirac neutrino spectra where there exists hierarchies among the new squared mass splitting. In particular, for the case which involves large $\delta m_{3}^{2}$ in comparison to other pseudo-Dirac mass splittings, a conservative upper bound of order $\delta m_{3}^{2}\lesssim 10^{-12}\,\textrm{eV}^{2}$ is obtained considering the size of our galaxy and assuming $\pi$-decay at neutrino sources with IceCube-Gen2 experiment. From eq. (35), if extragalatic neutrino sources are determined, the bound will greatly improve, up to five orders of magnitude if the sources are of the distance of Gpc. The benefit of using flavor ratios is its insensitivity to absolute neutrino flux uncertainty while the main caveat is that if all pseudo-Dirac mass splittings are of the same order, one will obtain an overall reduction of flux, making this method futile. As more astrophysical neutrino sources are identified, one can combine the flavor ratio method together with direct measurements to derive the best constraints on pseudo-Dirac neutrinos, exploring completely the window of pseudo-Dirac leptogenesis proposed in ref. Fong et al. (2021).
@@ Acknowledgments
YP acknowledges the support by Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) Contract No. 2023/10734-3 and 2023/01467-1. CSF acknowledges the support by FAPESP Contracts No. 2019/11197-6, 2022/00404-3 and Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under Contract No. 304917/2023-0.
@@
Note Added: A recent work Dev et al. (2024) has also considered effects of pseudo-Dirac neutrinos on astrophysical neutrino flavor ratios and showed that if there is an overdensity of cosmic neutrino background, matter effect can be relevant.
Appendix A $\mu$-damped production combined with pseudo-Dirac scenario
@@
In this section we show the results for the combination of $\mu$-damped production with the pseudo-Dirac scenario. We highlight that only in Case 12 the pseudo-Dirac scenario appears as clearly distinct from standard scenarios, although it will still be hard to discriminate even after 10 years of IceCube-Gen2 operation. As for the pseudo-Dirac scnearios involving $\nu_{3}$ mass eigenstates (Case 3, Case 13 and Case 23) assuming $\mu$-damped sources, it can result in confusion with the standard scenario assuming $\pi$-decay sources.
@@
Figure 3: Same as fig. 1, but with the assumption that pseudo-Dirac neutrinos are produced via the $\mu$-damped channel and corresponding projection of IceCube-Gen2 is taken from Abbasi et al. (2023b).
@@
Figure 4: Same as fig. 2, but with the assumption that pseudo-Dirac neutrinos are produced via the $\mu$-damped channeland corresponding projection of IceCube-Gen2 is taken from Abbasi et al. (2023b).

### v3

@@ V Conclusions
In this work, we have shown that using astrophysical neutrino flavor ratios, it is possible to distinguish between six possible cases of pseudo-Dirac neutrino spectra where there exists hierarchies among the new squared mass splitting. Our main finding is that the most sensitive cases turn out to involve the pseudo-Dirac mass splitting of the third mass eigenstates $\delta m_{3}^{2}$ which currently has the weakest constraint of the order $10^{-5}\,{\rm eV}^{2}$. For the cases $\delta m_{1,2}^{2}=0$ or $\delta m_{1}^{2}=0$, a conservative upper bound of order $\delta m_{3}^{2}\lesssim 10^{-12}\,\textrm{eV}^{2}$ can be obtained considering astrophysical neutrinos of energies $E\sim 100$ TeV coming from a distance greater than the size of our galaxy and assuming $\pi$-decay at neutrino sources with IceCube-Gen2 experiment. If nature is kind to us, we might detect the first positive evidence of pseudo-Dirac neutrino with $\delta m_{3}^{2}\sim 10^{-12}\,\textrm{eV}^{2}$ with astrophysical neutrino flavor ratio. From eq. (35), if extragalatic neutrino sources are determined, the bound will greatly improve, up to five orders of magnitude if the sources are of the distance of Gpc. The benefit of using flavor ratios is its insensitivity to absolute neutrino flux uncertainty while the main caveat is that if all pseudo-Dirac mass splittings are of the same order, one will obtain an overall reduction of flux, making this method futile. As more astrophysical neutrino sources are identified, one can combine the flavor ratio method together with direct measurements to derive the best constraints on pseudo-Dirac neutrinos, exploring completely the window of pseudo-Dirac leptogenesis proposed in ref. Fong et al. (2021).
@@ Acknowledgments
Y.P. acknowledges support by Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) Contracts No. 2023/10734-3 and No. 2023/01467-1, and by the National Council for Scientific and Technological Development (CNPq) Grant No. 151168/2023-7. C.S.F. acknowledges support by FAPESP Contracts No. 2019/11197-6 and No. 2022/00404-3 and Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under Contracts No. 407149/2021-0 and No. 304917/2023-0.
@@
Note Added: A recent work Dev et al. (2024) has also considered effects of pseudo-Dirac neutrinos on astrophysical neutrino flavor ratios and showed that if there is an overdensity of cosmic neutrino background by a factor of $10^{4}$ around the Earth, matter effect can be relevant. Since gravitational effect can only give rise to order of one enhancement, further new physics is required to achieve this large enhancement. To be conservative, we will assume no such large overdensity and ignore the neutrino matter effect.
Appendix A Appendix
$\mu$-damped production and the combination with $\pi$-decay and neutron decay production
@@ Appendix A Appendix
We first show the results for the combination of $\mu$-damped production with the pseudo-Dirac scenario. We highlight that only in Case 12 the pseudo-Dirac scenario appears as clearly distinct from standard scenarios, although it will still be hard to discriminate even after 10 years of IceCube-Gen2 operation. As for the pseudo-Dirac scenarios involving $\nu_{3}$ mass eigenstates (Case 3, Case 13 and Case 23) assuming $\mu$-damped sources, it can result in confusion with the standard scenario assuming $\pi$-decay sources, see figs. 3 and 4. In Table 3, we list the best-fit flavor fractions for each of the cases illustrated in fig. 3 and fig. 4.
At the Source
[…]
At the Earth
Scenario ($f_{e,\oplus}^{\text{det}}$, $f_{\mu,\oplus}^{\text{det}}$, $f_{\tau,\oplus}^{\text{det}}$)
Std. oscillations (0.18, 0.45, 0.37)
Case 1 (0.16, 0.46, 0.38)
Case 2 (0.15, 0.47, 0.38)
Case 3 (0.24, 0.41, 0.35)
Case 12 (0.12, 0.49, 0.39)
Case 13 (0.24, 0.42, 0.34)
Case 23 (0.23, 0.41, 0.36)
Table 3: Flavor composition at Earth for different scenarios using the best fit oscillation parameters of ref. de Salas et al. (2021), assuming muon damped at the source.
In fig. 5, we show ternary plots for the allowed regions in both the standard and pseudo-Dirac scenarios for the case in which $\pi$-decay and $\mu$-damped contributions, in varying proportions, are simultaneously present. To show how this can affect the two most interesting cases (Cases 3 and 23), we present the analysis for them. The result is that corresponding regions shown in figs. 1-4 merge significantly. The blue region corresponds to a mixture of $\mu$-damped and $\pi$-decay contributions in varying proportions, assuming the standard scenario. It largely overlaps with the red region, which represents the same mixture but in the presence of pseudo-Dirac neutrinos (the purple or dark red region indicates the overlap between the blue and light red areas).
In fig. 6, we add to the set of possible production processes the neutron decay, which is also a benchmark scenario often quoted in the literature. Neutron decay produces neutrinos with source flavor ratios $(1,0,0)_{S}$, which evolve into an approximate composition of $(0.55,0.17,0.28)_{\oplus}$ at Earth, thereby shifting the allowed region in the flavor triangle toward its lower-right corner in both standard and pseudo-Dirac scenarios Song et al. (2021).
@@
Figure 3: Same as fig. 1 in the main text, but with the assumption that pseudo-Dirac neutrinos are produced via the $\mu$-damped channel and corresponding projection of IceCube-Gen2 is taken from Abbasi et al. (2023b).
@@
Figure 4: Same as fig. 2 in the main text, but with the assumption that pseudo-Dirac neutrinos are produced via the $\mu$-damped channel and corresponding projection of IceCube-Gen2 is taken from Abbasi et al. (2023b).
Figure 5: Expected regions for the flavor composition of high-energy astrophysical neutrinos in the standard scenario (blue) and the pseudo-Dirac scenario (red). In both cases, we vary the oscillation parameters and consider two production mechanisms (pion decay and muon-damped) of varying proportions to obtain the allowed regions. Note that the standard and pseudo-Dirac scenarios largely overlap, except for the lower portion of the red region (originating from pion-decay sources emitting pseudo-Dirac neutrinos; see figs. 1 and 2), which could, in principle, be distinguished by future experiments such as IceCube-Gen2.
Figure 6: Same as Fig. 5, but in addition to pion decay and muon-damped production, we include neutron decay as a possible production mechanism. Neutron decay leads to source flavor ratios of the form $(1,0,0)_{S}$ and corresponds to flavor ratios at Earth approximately $(0.55,0.17,0.28)_{\oplus}$, which pushes the allowed region toward the lower-right vertex of the flavor triangle Song et al. (2021).